# State-Space Modeling in Long Sequence Processing: A Survey on Recurrence in the Transformer Era

### v1

@@ 2 Recurrent Models in Sequence Processing
where $A^{\prime}$ is the matrix transpose, $\text{diag}(\cdot)$ converts a vector into a diagonal matrix, and $\hat{\sigma}$ is the derivative of the activation function $\sigma$ in Eq. (2), which is applied in an element-wise fashion to its input vector. Unless the partial terms $\frac{\partial x_{s}}{\partial x_{s-1}}$ are close to 1, the product in Eq. (6) could explode or vanish [104]. In details, in the simplified case of a linear model (i.e., replacing $\sigma$ with the identity function) the power iteration method helps in deriving tight boundaries. In particular, a sufficient condition for long term information to vanish is that of having the largest eigenvalue of the recurrent weight matrix $A$ smaller than one, i.e. $\lambda<1$. Conversely, $\lambda>1$ is a necessary condition for gradients to explode (see [219] for further details). When gradients undergo vanishing during their backward propagation through time, the critical credit assignment aspect of backpropagation is compromised. In particular, the information about minor state changes in the distant past loses its ability to influence future states. Conversely, when gradients explode, gradient-based optimization algorithms encounter substantial difficulties in traversing the cost surface. This is primarily due to the fact that gradient-based optimization relies on the assumption that small parameter adjustments yield small changes in the objective function. As the number of time steps considered in the sequence of states increases, the amplifying or diminishing effects associated with the state-to-state transformations at individual time steps can grow exponentially, making it difficult for these models to discern long-range dependencies in the data.
@@
Properties of Recurrent Models.   The sum-product of $L$ terms in Eq. 5 results in $\mathcal{O}(L^{2})$ scaling for each processed sequence. In practice, efficient gradient propagation schemes have been devised in order to result in a cost linear in length of the sequence [130, 84]. In terms of memory consumption, BPTT requires to store all the $L$ intermediate states, resulting in a $\mathcal{O}(L)$ overhead (see also [84] for alternative methods to achieve better memory efficient BPTT). Being RNN state-full models, it is not possible to parallelize their inference/training procedure over the time instants in to which the components of the sequence belong. [^3] See Section 4 for special cases that can be parallelized. Computing the $t$-th state is conditioned on the current input $u_{t}$ and the preceding state $x_{t-1}$, that summarizes all the past knowledge and behaves as an information bottleneck. In other words, RNNs are causal models, since they can only leverage past information available in the input sequence. Indeed, to produce an output $y_{t+1}$, the model can only access past information up to $u_{t}$. A big advantage of this computational scheme is that, during inference, RNNs only require constant computation/storage per time step, i.e., $\mathcal{O}(1)$. From the theoretical point of view, RNNs have been proven to be able to simulate Universal Turing Machines [267], i.e., they can combine symbolic and sub-symbolic capabilities by running algorithms on a neural substrate. In this direction, some works introduced the unrealistic assumption of neurons with unbounded precision that equals the number of symbols used in the Turing tape [268, 267]. Chung et al. [42] relaxed such assumptions by leveraging a dynamically growing memory module made of neurons of fixed precision. Additionally, RNNs are Universal Approximators [257], i.e., they are able to approximate any open dynamical system with an arbitrary accuracy (see [160] and references therein for further Universal Approximation results for both discrete and continuous RNNs).
@@
Other Surveys on Recurrent Models. Due to the large popularity of RNNs, there are several books [81], surveys and reviews that describe their properties and architectural advances. Detailed overviews were included in recent surveys [169, 253], up to models presented in 2015-2017, respectively. An in-depth analysis on gating-based architectures has been given in [310]. Bianchi et al. [21] start from a general overview on RNNs and learning algorithms to later focus on short-term load forecasting. The BPTT algorithm and its biological plausibility are investigated in [163]. An important branch of research looks towards the discovery of online learning algorithms in the context of RNNs, and [184] proposes an interesting framework that summarizes all the recent findings. There are also survey that focus on the application of RNNs in confined areas, such as continual [49, 62] and online learning [216].
@@
Recent Findings.   The usage of Recurrent Models for sequence processing became less popular since 2017, due to the already enstablished popularity of Convolutional models [13] and, more importantly, due to the growing ubiquity of Transformer-based solutions [291], pushed by the ease of parallelizing training over the time instants on which the input sequences are defined. However processing long sequences is hard and expensive with such architectures, fostering the need of novel paths to achieve efficient architectures [283]. As a result, the scientific literature recently experienced the resurgence of interest in RNNs [134, 213]. Transformers’ quadratic complexity on the sequence length has pushed towards the discovery of novel methods to optimize inference [283, 166], and many solutions based on stateful recurrent computations have been introduced [134, 222, 227, 277]. At the same time, the need to preserve extremely long-term dependencies on sequences has led to the adoption of the so-called (deep) State-Space Models (SSMs) [86, 271]. This research field originated from the existing knowledge on RNNs [293], but then it took its own way. It has been very recently formally re-connected to RNNs by a careful architectural design [213, 57]. Of course, this is not surprising, since, as anticipated, RNNs are indeed instances of SSMs.
@@ 3 Transformers Embracing Recurrent Models
Figure 5: Overview of the organization of Section 3. We report the names of the subsections (blue boxes) and the covered categories of models (yellow). The names of some representative methods are also indicated for each category (gray).
Transformers [291] emerged as a disruptive alternative to Recurrent Models, promoting the idea of going beyond stateful models which process the sequence one element at a time, motivated by the “attention is all you need” motto. Transformers are able to handle the elements of input sequence [283] in parallel and to capture local and long-range dependencies. These properties are due to their self-attention-based computational scheme, that compares all the possible pairs of input patterns constituting the sequence. Here we will start from a brief analysis aimed at summarizing the properties of self-attention and, consequently, we will underline its drawbacks. Finally, we will delve into the latest findings on self-attention variants and alternatives that, surprisingly, can be seen though the lens of Recurrent Models. Figure 5 showcases the organization of this section.
Transformers and Self-Attention.   Given the input sequence $(u_{1},u_{2},\dots,u_{L})$ with $u_{t}\in\mathbb{R}^{d_{\text{in}}}$, $t=1,\ldots L$, Transformers implements a sequence-to-sequence function $\mathrm{tr}\colon\mathbb{R}^{L\times{d_{\text{in}}}}\to\mathbb{R}^{L\times{d_{y}}}$, that yields $(y_{1},y_{2},\dots,y_{L})$. The whole function $\mathrm{tr}$ is based on a self-attention procedure, followed by a feed-forward network, and it is commonly applied multiple times in a multi-layer fashion. [^4] We purposely avoid to describe these components and others, such as skip connections, layer normalization, multi-head attention. We also do not mention the distinction between Transformer encoders and decoders, since they do not introduce useful information for the point we make on this survey. Please refer to [283] and references therein for detailed descriptions of the Transformers architecture. Self-attention acts across the temporal dimension of the input sequence, evaluating pairwise interactions between the input components, regardless of their position in the sequence, [^5] Positional embeddings are commonly exploited to make Transformers position-dependent. and it is evaluated (in parallel) on each element of the input sequence. When evaluated on a generic $u_{t}$, it returns a sum over $(u_{1},u_{2},\dots,u_{L})$, weighed by scores that depend on the similarities between $u_{t}$ and the $L$ elements of the input sequence. Formally,
@@
where $q_{\cdot}$, $k_{\cdot}$, $v_{\cdot}$ are the so-called queries, keys and values, respectively, computed by projecting (three linear projections) the elements of the input sequence using three trainable matrices $W_{q},W_{k}\in\mathbb{R}^{{d_{\text{k}}}\times{d_{\text{in}}}}$ and $W_{v}\in\mathbb{R}^{{d_{y}}\times{d_{\text{in}}}}$, where ${d_{\text{k}}}$ is the keys and query size while ${d_{y}}$ is the self-attention output size. We have $q_{z}=W_{q}u_{z}$, $k_{z}=W_{k}u_{z}$ and $v_{z}=W_{v}u_{z}$, for $z=1,\ldots,L$. The similarity function $\text{sim}\left(q_{\cdot},k_{\cdot}\right)$ returns a positive score which is higher in case of larger similarity between its arguments. The canonical choice for the self-attention function is the softmax dot-product (multiplicative) attention, where the similarity score is computed as the exponential of the dot product between a query and a key, $\text{sim}\left(q_{\cdot},k_{\cdot}\right)=\exp\left({q_{\cdot}^{\prime}k_{\cdot}/\sqrt{d_{k}}}\right)$, where is the transpose operator and all mono-dimensional arrays are intended to be column vectors (here and in the rest of the paper). Before going into further details, we introduce the matrix notation that we will use throughout this section, to simplify several descriptions. Matrix $U$ is composed of $L$ rows, each of them storing the transpose of $u_{t}$, for $t=1,\ldots,L$. Similarly, for all the already introduced symbols that are associated to the $L$ elements of the input sequence, their uppercase versions (with no subscripts) indicate matrices with $L$ rows (e.g., in the case of $y_{t}$, $q_{t}$, $k_{t}$, $v_{t}$, $t=1,\ldots,L$, we have matrices $Y$, $Q$, $K$, $V$, respectively). We can rewrite Eq. 7 in matrix notation as follows,
@@
where $\mathrm{softmax}(A)$ is the softmax function that operates on each row of $A$. The above formulation is usually referred to as the parallel form of attention, since it computes the Transformer outputs for all the time instants of the sequence (i.e., the full matrix $Y$) in parallel given the full input matrix $U$, i.e. all the sequence must be available in advance. This is one of the main advantages of Transformers: during training, which is usually performed leveraging the full input sequence, computations can be efficiently parallelized over the sequence dimension – hence the term parallel training.
@@
Causal Attention.   Differently from the aforementioned scenario where the full sequence is available in advance, there exists settings where future information in the sequence cannot be accessed, thus requiring to implement a form of causal attention. For instance, Transformers can work in an autoregressive setting, where the input sequence is progressively populated in an iterative manner (adding an element predicted by the Transformer itself at each time step, e.g., during inference while decoding). The standard self-attention of Eq. (7) is not causal, given that future position $j>i$ can influence the current one. However, autoregressive Transformers can be obtained by causal masking, i.e., masking the attention computation such that the $i$-th position can only be influenced by precedings ones. In this causal setting, the parallel form of Eq. (8) can be rewritten as $Y=\mathrm{softmax}\left(\frac{QK^{\prime}}{\sqrt{d_{k}}}\odot{M}\right)V$, where $M\in\mathbb{R}^{L\times L}$ is a mask that prevents the model from attending to future tokens, i.e., $M_{ij}=1$ if $i\geq j$ and $M_{ij}=-\infty$ if $i<j$. During inference, causal Transformers can be interpreted as exploiting a recurrent form, which is obtained by considering only the time steps up to the current one in Eq. (7), i.e., $y_{t}=\sum_{i=1}^{t}\frac{\text{sim}\left(q_{t},k_{i}\right)}{\sum_{j=1}^{t}\text{sim}\left(q_{t},k_{j}\right)}v_{i},\ t=1,\ldots,L.$ Notice that, through this causal mechanism, attention is performed over sets of keys and values that grow over time: to compute $y_{t-1}$, the sequences $(k_{i})_{i=1}^{t-1}$ and $(v_{i})_{i=1}^{t-1}$ are needed; to compute $y_{t}$, the additional key and value $k_{t}$ and $v_{t}$ must be considered in the computation as well, thus expanding the aforementioned sequences of keys and values (over time)—also referred to as KV-cache [228].
@@
Computational Cost and Scaling Issues.   As stated above, one of the key contribution of Transformers architecture is that inference can be parallelized over the temporal dimension [283], if the full sequence is available in advance, going beyond the intrinsic limitations of stateful Recurrent Models. In fact, all the matrix-by-matrix products in Eq. 8 can be computed as $L$ independent vector-by-matrix products. Thus, there is evident room for parallel implementations, given the full input $U$. Despite the significant benefits brought by parallelization, Transformers are characterized by serious computational and memory costs. In fact, differently from Recurrent Models that keep track of the prior context via an explicit state with fixed size, Transformers self-attention has a “global” receptive field that directly attends the whole sequence at inference. As a result, the main bottleneck is constituted by the self-attention, which scales quadratically with the sequence length, $\mathcal{O}(L^{2})$. Indeed, from the definition in Eq. 8: computing $QK^{\prime}/\sqrt{d_{k}}$ takes $\mathcal{O}(L^{2}{d_{\text{k}}})$; the $\mathrm{softmax}$ function involves elementwise exponentiation (and sum) of its matrix argument, taking $\mathcal{O}(L^{2})$ time, and division of each element with the corresponding row sum, which takes $\mathcal{O}(L^{2})$; multiplication of $\mathrm{softmax}(\cdot)$ and $V$ takes $\mathcal{O}\left(L^{2}\text{max}({d_{\text{k}}},{d_{y}})\right)$ time [136]; overall, we inference is then $\mathcal{O}(L^{2})$ with respect to the sequence length. In the aforementioned autoregressive setting, where the input sequence is progressively populated and keys and values grow over time in the KV-cache, the computational cost of autoregressive inference is $\mathcal{O}(L_{t})$, being $L_{t}$ the length of the accumulated sequence at time $t$. All the aforementioned complexities makes it harder to apply Transformers when scaling up the input sequence length. This issue forced practioners to artificially limit the sequence size, considering a custom context window length, hindering the model capability of capturing long-term dependencies. Investigations on such limitations spurred a number of efforts aimed at improving Transformers scalability [283], with the final goal of retaining their performances, parallel computations, and merge them with efficient inference complexity. Among efforts aimed at optimizing the attention mechanism [17, 144, 40, 54] or to completely replace it [313], many recent works have been inspired by the advantages and peculiarities of Recurrent Models. We report an overview on the complexities of different variants of recent Transformer models in Table 3, which will be discussed in the following.

### v2

@@ 2 Recurrent Models in Sequence Processing
where $A^{\prime}$ is the matrix transpose, $\text{diag}(\cdot)$ converts a vector into a diagonal matrix, and $\hat{\sigma}$ is the derivative of the activation function $\sigma$ in Eq. (2), which is applied in an element-wise fashion to its input vector. Unless the partial terms $\frac{\partial x_{s}}{\partial x_{s-1}}$ are close to 1, the product in Eq. (6) could explode or vanish [57]. In details, in the simplified case of a linear model (i.e., replacing $\sigma$ with the identity function) the power iteration method helps in deriving tight boundaries. In particular, a sufficient condition for long term information to vanish is that of having the largest eigenvalue of the recurrent weight matrix $A$ smaller than one, i.e. $\lambda<1$. Conversely, $\lambda>1$ is a necessary condition for gradients to explode (see [56] for further details). When gradients undergo vanishing during their backward propagation through time, the critical credit assignment aspect of backpropagation is compromised. In particular, the information about minor state changes in the distant past loses the capacity to influence future states. Conversely, when gradients explode, gradient-based optimization algorithms encounter substantial difficulties in traversing the cost surface. This is primarily due to the assumption that gradient-based optimization relies on that small parameter adjustments yield small changes in the objective function. As the number of time steps considered in the sequence of states increases, the amplifying or diminishing effects associated with the state-to-state transformations at individual time steps can grow exponentially, making it difficult for these models to capture long-range dependencies in the data.
@@
Properties of Recurrent Models.   The sum-product of $L$ terms in Eq. 5 results in $\mathcal{O}(L^{2})$ scaling for each processed sequence. In practice, efficient gradient propagation schemes have been devised to achieve a cost linear in the length of the sequence [58, 59]. In terms of memory consumption, BPTT requires storing all the $L$ intermediate states, resulting in a $\mathcal{O}(L)$ overhead (see also [59] for alternative methods to achieve more memory-efficient BPTT). Since RNN are stateful models, it is not possible to parallelize their inference/training procedure over the time instants to which the components of the sequence belong. [^4] See Section 4 for special cases that can be parallelized. Computing the $t$-th state is conditioned on the current input $u_{t}$ and the preceding state $x_{t-1}$, which summarizes all the past knowledge and behaves as an information bottleneck. In other words, RNNs are causal models, since they can only leverage past information available in the input sequence. Indeed, to produce an output $y_{t+1}$, the model can only access past information up to $u_{t}$. A major advantage of this computational scheme is that, during inference, RNNs only require constant computational and storage resources per time step, i.e., $\mathcal{O}(1)$. From the theoretical point of view, RNNs have been proven to be able to simulate Universal Turing Machines [60], i.e., they can combine symbolic and sub-symbolic capabilities by running algorithms on a neural substrate. In this direction, some works introduced the unrealistic assumption of neurons with unbounded precision that equals the number of symbols used in the Turing tape [61, 60]. Chung et al. [62] relaxed such assumptions by leveraging a dynamically growing memory module made of neurons of fixed precision. Additionally, RNNs are Universal Approximators [63], i.e., they are able to approximate any open dynamical system with an arbitrary accuracy (see [64] and references therein for further Universal Approximation results for both discrete and continuous RNNs). We further mention several studies on the way recurrent models (with trained and untrained state transition function) handle memorization, such as those that fall in the context of Echo State Networks [65]. The reader can refer to Section 6 for further details on the literature on this topic.
@@
Other Surveys on Recurrent Models. Due to the large popularity of RNNs, there are several books [66], surveys, and reviews that describe their properties and architectural advances. Detailed overviews were included in recent surveys [5, 18], covering models up to 2015–2017, respectively. An in-depth analysis of gating-based architectures has been given in [67]. Bianchi et al. [10] provide a general overview on RNNs and learning algorithms before focusing on short-term load forecasting. The BPTT algorithm and its biological plausibility are investigated in [68]. An important branch of research looks towards the discovery of online learning algorithms in the context of RNNs, and [45] proposes an interesting framework that summarizes recent findings. There are also surveys that focus on the application of RNNs in confined areas, such as continual [69, 9] and online learning [70].
@@
Recent Findings.   The usage of Recurrent Models for sequence processing became less popular since 2017, due to the already established popularity of Convolutional models [33] and, more importantly, due to the growing ubiquity of Transformer-based solutions [26], driven by the ease of parallelizing training over the time instants on which the input sequences are defined. However,processing long sequences is hard and expensive with such architectures, fostering the need for novel paths to achieve efficient architectures [71]. As a result, the scientific literature has recently experienced a resurgence of interest in RNNs [72, 43]. Transformers’ quadratic complexity with respect to the sequence length has pushed towards the discovery of novel methods to optimize inference [71, 73], and many solutions based on stateful recurrent computations have been introduced [72, 38, 74, 75]. At the same time, the need to preserve extremely long-term dependencies on sequences has led to the adoption of the so-called (deep) State-Space Models (SSMs) [39, 42]. This research field originated from the existing knowledge on RNNs [30], but then it took its own direction. It has been very recently formally reconnected to RNNs by a careful architectural design [43, 44]. This is not surprising, since, as anticipated, RNNs are indeed instances of SSMs. In this section, we established the foundational dichotomy between stateless sequence processors and stateful recurrent models, formalised the Elman-type RNN, and analysed BPTT together with its vanishing/exploding gradient pathology. The discussion highlighted RNNs’ theoretical universality, their $\mathcal{O}(1)$ inference cost, and their recent resurgence prompted by the scaling limits of Transformers. These insights set the stage for hybrid approaches that attempt to combine recurrent efficiency with modern attention mechanisms.
The classical RNN framework outlined above clarifies both its expressive power and its well-known training bottlenecks (vanishing/exploding gradients and limited parallelism). These limitations have recently motivated a wave of research that injects recurrence into the dominant Transformer family, seeking the best of both worlds—parallel training and explicit long-range memory. The next section surveys this convergence and shows how “attention meets recurrence”.
@@ 3 Transformers Embracing Recurrent Models
Figure 5: Conceptual overview of the organization of Section 3,which categorizes Transformer-based models according to architectural and functional aspects. We report the names of the subsections (blue boxes) and the covered categories of models (yellow). The names of some representative methods are also indicated for each category (gray).
Transformers [26] emerged as a disruptive alternative to Recurrent Models, promoting the idea of going beyond stateful models that process the sequence one element at a time, motivated by the “attention is all you need” motto. Transformers are able to handle the elements of an input sequence [71] in parallel and to capture both local and long-range dependencies. These properties are due to their self-attention-based computational scheme, which compares all the possible pairs of input patterns constituting the sequence. In the following, we present a brief analysis aimed at summarizing the properties of self-attention and subsequently underline its drawbacks. Finally, we will delve into the latest findings on self-attention variants and alternatives that, perhaps unexpectedly, can be seen through the lens of Recurrent Models. Figure 5 showcases the organization of this section.
Transformers and Self-Attention.   Given the input sequence $(u_{1},u_{2},\dots,u_{L})$ with $u_{t}\in\mathbb{R}^{d_{\text{in}}}$, $t=1,\ldots,^{L}$, Transformers implement a sequence-to-sequence function $\mathrm{tr}\colon\mathbb{R}^{L\times{d_{\text{in}}}}\to\mathbb{R}^{L\times{d_{y}}}$, that yields $(y_{1},y_{2},\dots,y_{L})$. The whole function $\mathrm{tr}$ is based on a self-attention procedure, followed by a feed-forward network, and it is commonly applied multiple times in a multi-layer fashion. [^5] We purposely avoid describing these components and others, such as skip connections, layer normalization, and multi-head attention. We also do not mention the distinction between Transformer encoders and decoders, since they do not introduce useful information for the point we make. Please refer to [71] and references therein for detailed descriptions of the Transformer architecture. Self-attention acts across the temporal dimension of the input sequence, evaluating pairwise interactions between the input components, regardless of their positions in the sequence, [^6] Positional embeddings are commonly exploited to make Transformers position-dependent. and it is evaluated (in parallel) on each element of the input sequence. When evaluated on a generic $u_{t}$, it returns a sum over $(u_{1},u_{2},\dots,u_{L})$, weighed by scores that depend on the similarities between $u_{t}$ and the $L$ elements of the input sequence. Formally,
@@
where $q_{\cdot}$, $k_{\cdot}$, $v_{\cdot}$ are the so-called queries, keys and values, respectively, computed by projecting (three linear projections) the elements of the input sequence using three trainable matrices $W_{q},W_{k}\in\mathbb{R}^{{d_{\text{k}}}\times{d_{\text{in}}}}$ and $W_{v}\in\mathbb{R}^{{d_{y}}\times{d_{\text{in}}}}$, where $d_{\text{k}}$} is the keys and query size while $d_{y}$} is the self-attention output size. We have $q_{z}=W_{q}u_{z}$, $k_{z}=W_{k}u_{z}$ and $v_{z}=W_{v}u_{z}$, for $z=1,\ldots,L$. The similarity function $\text{sim}\left(q_{\cdot},k_{\cdot}\right)$ returns a positive score that increases with the similarity between its arguments. The canonical choice for the self-attention function is the softmax dot-product (multiplicative) attention, where the similarity score is computed as the exponential of the dot product between a query and a key, $\text{sim}\left(q_{\cdot},k_{\cdot}\right)=\exp\left({q_{\cdot}^{\prime}k_{\cdot}/\sqrt{d_{k}}}\right)$, where is the transpose operator and all mono-dimensional arrays are assumed to be column vectors (here and in the rest of the paper). Before going into further details, we introduce the matrix notation that we will use throughout this section to simplify the descriptions. Matrix $U$ is composed of $L$ rows, each storing the transpose of $u_{t}$, for $t=1,\ldots,L$. Similarly, for all the already introduced symbols that are associated with the $L$ elements of the input sequence, their uppercase versions (with no subscripts) indicate matrices with $L$ rows (e.g., in the case of $y_{t}$, $q_{t}$, $k_{t}$, $v_{t}$, $t=1,\ldots,L$, we have matrices $Y$, $Q$, $K$, $V$, respectively). We can rewrite Eq. 7 in matrix notation as follows,
@@
where $\mathrm{softmax}(A)$ is the softmax function that operates on each row of $A$. The above formulation is usually referred to as the parallel form of attention, since it computes the Transformer outputs for all the time instants of the sequence (i.e., the full matrix $Y$) in parallel given the full input matrix $U$, meaning that all the sequence must be available in advance. This is one of the main advantages of Transformers: during training, which is usually performed using the full input sequence, computations can be efficiently parallelized over the sequence dimension—hence the term parallel training.
@@
Causal Attention.   Unlike the aforementioned scenario where the full sequence is available in advance, there are settings where future information in the sequence cannot be accessed, thus requiring the implementation of a form of causal attention. For instance, Transformers can work in an autoregressive setting, where the input sequence is progressively populated in an iterative manner (adding an element predicted by the Transformer itself at each time step, e.g., during inference while decoding). The standard self-attention of Eq. (7) is not causal, given that future positions $j>i$ can influence the current one. However, autoregressive Transformers can be obtained by causal masking, i.e., masking the attention computation such that the $i$-th position can only be influenced by preceding ones. In this causal setting, the parallel form of Eq. (8) can be rewritten as $Y=\mathrm{softmax}\left(\frac{QK^{\prime}}{\sqrt{d_{k}}}\odot{M}\right)V$, where $M\in\mathbb{R}^{L\times L}$ is a mask that prevents the model from attending to future tokens, i.e., $M_{ij}=1$ if $i\geq j$ and $M_{ij}=$-$\infty$ if $i<j$. During inference, causal Transformers can be interpreted as exploiting a recurrent form, which is obtained by considering only the time steps up to the current one in Eq. (7), i.e., $y_{t}=\sum_{i=1}^{t}\frac{\text{sim}\left(q_{t},k_{i}\right)}{\sum_{j=1}^{t}\text{sim}\left(q_{t},k_{j}\right)}v_{i},\ t=1,\ldots,L.$ Note, that, through this causal mechanism, attention is performed over sets of keys and values that grow over time: to compute $y_{t-1}$, the sequences $(k_{i})_{i=1}^{t-1}$ and $(v_{i})_{i=1}^{t-1}$ are needed; to compute $y_{t}$, the additional key and value $k_{t}$ and $v_{t}$ must be considered in the computation as well, thus expanding the aforementioned sequences of keys and values over time also referred to as the KV-cache [76].
@@
Computational Cost and Scaling Issues.   As stated above, one of the key contributions of the Transformer architecture is that inference can be parallelized over the temporal dimension [71], if the full sequence is available in advance, thus overcoming the intrinsic limitations of stateful Recurrent Models. In fact, all the matrix-by-matrix products in Eq. 8 can be computed as $L$ independent vector-by-matrix products. Thus, there is evident room for parallel implementations, given the full input $U$. Despite the significant benefits brought by parallelization, Transformers are characterized by serious computational and memory costs. In fact, unlike Recurrent Models that keep track of the prior context via an explicit state with fixed size, Transformer self-attention has a “global” receptive field that directly attends the whole sequence at inference. As a result, the main bottleneck is constituted by the self-attention, which scales quadratically with the sequence length, $\mathcal{O}(L^{2})$. Indeed, from the definition in Eq. 8: computing $QK^{\prime}/\sqrt{d_{k}}$ takes $\mathcal{O}(L^{2}{d_{\text{k}}})$; the $\mathrm{softmax}$ function involves elementwise exponentiation (and sum) of its matrix argument, taking $\mathcal{O}(L^{2})$ time, and division of each element by the corresponding row sum, which takes $\mathcal{O}(L^{2})$; multiplication of $\mathrm{softmax}(\cdot)$ and $V$ takes $\mathcal{O}\left(L^{2}\text{max}({d_{\text{k}}},{d_{y}})\right)$ time [77]; overall, inference has $\mathcal{O}(L^{2})$ complexity in the sequence length.. In the aforementioned autoregressive setting, where the input sequence is progressively populated and keys and values grow over time in the KV-cache, the computational cost of autoregressive inference is $\mathcal{O}(L_{t})$, being $L_{t}$ the length of the accumulated sequence at time $t$. All the aforementioned complexities make it harder to apply Transformers when scaling up the input sequence length. This issue has forced practitioners to artificially limit the sequence size, considering a custom context window length, thus hindering the model’s capability of capturing long-term dependencies. Investigations on such limitations have spurred a number of efforts aimed at improving Transformers’ scalability [71], with the final goal of retaining their performance, parallel computations, and merging them with efficient inference complexity. Among efforts aimed at optimizing the attention mechanism [78, 79, 80, 37] or completely replacing it [81], many recent works have been inspired by the advantages and peculiarities of Recurrent Models. We report an overview of the complexities of different variants of recent Transformer models in Table 4, which will be discussed in the following.